Order: Heparin 20,000 units in​ 1,000 mL​ 0.9% NaCl​ stat.,infuse at 25​ mL/h.

Calculate the units per hour and determine if the dose is within the normal daily range​ (20,000-40,000 units in 24​h).

0.8​ units/h, and the dose is lower than the normal daily range.
500,000 units/h, and the dose is higher than the normal daily range.
25,000 units/h, and the dose is within the normal daily range.
500​ units/h, and the dose is lower than the normal daily range.
Answer #1

GIVEN,

Heparin- 20,000 units in 1000 ml 0.9% Na cl stat. infuse at 25 ml / hr.

1 ml = 20,000/1000 = 20 units

It is given that it must be infused at 25 ml/hr

= 25*20=500 units of heparin

that means 500 units/hr

it is given that normal daily range is 20000-40000 in 24 hours

so, 500 * 24 = 12,000

so ,it is lower than the normal daily range

therefore , answer is D

500 units/hr lower than the normal daily range.